OpenStudy (anonymous):

I have no idea how to do this problem. I know how to do others, but the absolute value is throwing me off. Any help is appreciated

OpenStudy (anonymous):

Absolute value is the number of units the number is away from 0 on a number line, so the absolute value of a number is always positive.

OpenStudy (anonymous):

\[\int\limits_{-\infty}^{\infty}e ^{-\left[ x \right]}\]

OpenStudy (anonymous):

Evaluate the improper integral

OpenStudy (anonymous):

I forgot the dx

OpenStudy (anonymous):

I know how to fix the infinite bounds issue, but the absolute value exponent is the issue for me

OpenStudy (anonymous):

@Zarkon If you could help I would really appreciate it

OpenStudy (anonymous):

@jdoe0001

OpenStudy (irishboy123):